# Compress, Gather, and Recompute: REFORMing Long-Context Processing in Transformers

### v1

@@ 4.6 Efficiency Analysis
Table 7: Efficiency Analysis. We compare the peak memory usage and inference time required for generating 10 tokens conditioned on 256k inputs. All measurements are made with the Mistral-NeMo-Instruct-2407 model on a single H100 GPU, and are averaged over 10 samples. The best values are highlighted in bold.
Inference Time (sec.) Peak Memory (GB) StreamingLLM 36.58 37.34 H2O 41.33 37.85 TOVA 39.46 37.06 InfiniPot 40.90 37.06 InfLLM 129.14 51.62 REFORM (Ours) 27.24 35.00
We outline the results in Table 7. InfLLM suffers from high inference time due to frequent memory transfer between CPU and GPU despite requiring large memory to store the cache. Recurrent methods offer faster inference at lower memory costs, enjoying the benefits of using a fixed-size KV cache. Interestingly, our approach shows lower latency and memory requirements compared to the recurrent baselines, despite having to forward the gathered inputs again and storing token-level context embeddings in memory. This is mainly due to the savings from early exit, which saves computation as well as memory by removing the need to keep the KV cache for the upper Transformer layers.
5 Conclusion
In this work, we introduced REFORM, a novel inference framework for efficient long-context processing. REFORM incrementally processes input chunks while maintaining a compressed KV cache, extracting key QKV states to construct cross-layer context embeddings. An early-exit strategy enhances efficiency, and a similarity-based selection mechanism identifies and gathers essential tokens for KV cache recomputation, ensuring accurate long-range information retention. REFORM outperforms existing methods across long-context benchmarks while reducing inference time and memory usage. Furthermore, its modality-agnostic design makes it applicable to a wide range of use cases including multi-modal applications.
Future work. In this paper, we directly adopt the token eviction criteria proposed by H2O (Zhang et al., 2023) as the compression component of our framework. As a future direction, we intend to investigate more sophisticated compression approaches that are better tailored for constructing the context embeddings. Furthermore, applications to more diverse data modalities such as audio and video is another promising direction to explore.
Impact Statement
As the demand for processing vast amounts of information grows across various fields, the ability to handle extremely long-context inputs efficiently becomes increasingly vital. REFORM presents a significant advancement in long-context processing, enabling large language models to reason over extremely long sequences without prohibitive computational costs.
By significantly improving efficiency, memory usage, and performance, REFORM makes advanced AI more accessible and sustainable. Its ability to retain and prioritize critical information while reducing computational overhead may support real-world applications, where processing extensive records with accuracy is essential.
Furthermore, REFORM contributes to AI democratization by lowering resource barriers, making high-performance long-context processing available to a broader range of users, including researchers, educators, and smaller enterprises. Its impact extends to environmental sustainability, as reducing computational demands translates to lower energy consumption in large-scale AI deployments.
By bridging the gap between efficiency and performance, REFORM not only advances the technical landscape of AI but also fosters a future where AI-driven insights are more inclusive, responsible, and impactful across diverse societal domains.
Appendix A Extended related works
Extending LLMs to handle extremely long inputs. To extend the context windows of Large Language Modles (LLMs) efficiently, various approaches have been proposed. A significant body of work focuses on modifying positional embeddings. These include scaling Rotary Positional Embeddings (RoPE) (Su et al., 2024) beyond the model’s context limit (Chen et al., 2023a; kaiokendev, 2023; bloc97, 2023; Peng et al., 2023), applying attention masks (Han et al., 2023), or adjusting the relative distances between tokens to fall within a predefined range (Su, 2023; Jin et al., 2024). Another line of research explores fine-tuning techniques to adapt models for longer contexts (Zhu et al., 2023; Chen et al., 2023b). While these methods enable models to handle extended inputs, they do not address the significant computational and memory costs introduced by the self-attention mechanism, limiting their practical utility for extremely long contexts.
Other approaches for efficient long context processing. Together with the recurrent KV cache compression approaches, a large volume of recent works focus on reducing the size of the KV cache to enable more efficient inference at long contexts. For example, SnapKV (Li et al., 2024a) proposes to forward the full input through the model, and then compress the cache by evicting tokens based on attention scores. While efficient at decoding-time, it requires the model to first process the full input, and therefore is not applicable to extremely long inputs that exceed the model’s pre-trained context window. Alternatively, HOMER (Song et al., 2024) proposes to use a hierarchical divide-and-conquer approach to combine the encoding and eviction process. Some works propose to further enhance KV cache compression by merging tokens instead of evicting them (Wang et al., 2024b; Dong et al., 2024), but their experiments also only consider inputs within the model’s context limit, and their extrapolation capabilities remain unknown. Some recent works propose another direction to keep the full cache only for some selected attention heads known as ‘retrieval heads’ (Tang et al., 2024; Xiao et al., 2024b), reducing the memory burden of preserving the full KV cache. Other works investigate quantization (Hooper et al., 2024; Kang et al., 2024) and low-rank cache compression (Singhania et al., 2024) to further reduce the memory requirements of the KV cache. However, these methods also cannot extrapolate to longer sequences beyond the model’s pre-trained context limit.
Appendix B Experimental Details
B.1 Evaluating Retrieval Heads and Embeddings
Dataset preparation. To evaluate the embeddings, we constructed a synthetic dataset based on multi-hop question answering. In this setup, we embedded documents from the HotPotQA dataset (Yang et al., 2018) at random positions within a long text corpus derived from the WikiText dataset (Merity et al., 2016). Each question was appended at the end of the context, and token-level labels were created, where tokens from the golden documents were marked as ground truth. All samples were designed to be 8k tokens long, which is within the context window of the Mistral-7B-Instruct-v0.2 model.
Embedding extraction. To simulate long-context scenarios where full attention computation is infeasible due to computational or memory constraints, we employed a recurrent chunk forwarding method based on H2O (Zhang et al., 2023), elaborated in Section 3.1. For attention, we compute the retrieval scores using the dot product between query states (Q) and the key states (K) without applying positional encoding. For all other embeddings, we compute the significance scores using cosine similarity between question embeddings and context embeddings, followed by max-pooling over question tokens. Additionally, retrieval scores for each context token were smoothed by mean-pooling with 20 neighboring tokens.
B.2 Multimodal Evaluations
B.3 Comparison against RAG
Experiment setup. We follow the setup in OP-RAG (Yu et al., 2024) for the RAG experiments, segmenting the inputs to 128-token chunks and preserving the order of the chunks instead of rearranging them according to the retrieval scores. We use Mistral-NeMo-Instruct-2407 as the base LLM. We use BM25 (Robertson et al., 1995) as the sparse retriever, and bge-large-en-v1.5 (Xiao et al., 2023b) as the dense retriever. For each sample, 8k tokens are retrieved in total, matching the KV size with our approach to ensure fair comparison.
B.4 Efficiency Measurements
Experiment setup. We measure the average inference time and peak memory usage for generating 10 tokens conditioned on 256k tokens. All measurements are made on a single H100 GPU, and we apply Flash Attention 2 (Dao, 2024) for all measurements. We further elaborate the experiment setup for InfLLM, as the inference speed and memory consumption can largely vary depending on the configuration. We use the default configuration provided in their GitHub repository, while modifying the number of retrieved blocks to keep 32k active tokens in the cache. The maximum number of blocks cached in GPU was set to be the twice as large as the number of retrieved blocks, following the convention in their official configuration file.
B.5 Embedding Construction and Similarity Search for REFORM
Embedding head selection. We construct the context embeddings by combining four QKV embeddings, where two heads are identified using the pattern matching dataset and the other two are identified using the multi-hop QA dataset. To balance between performance and efficiency gains, we select the top-performing heads from layers with depth under 70% for pattern matching heads. See Section C.2 for a more detailed discussion.
For Mistral-NeMo-Instruct-2407, the following heads are used:
  1. 1.
    Pattern Matching: Value heads 0 and 7 at layer 27
  2. 2.
    Multi-hop QA: Value head 5 at layer 19, query head 9 at layer 15
For Qwen2.5-7B-Instruct, the following heads are used:
  1. 1.
    Pattern Matching: Value head 3 at layer 7, key head 0 at layer 14
  2. 2.
    Multi-hop QA: Value head 3 at layer 14, value head 0 at layer 19
For Pixtral-12B-2409, the following heads are used:
  1. 1.
    Pattern Matching: Value heads 0 and 7 at layer 27
  2. 2.
    Multi-hop QA: Value head 3 at layer 10, value head 5 at layer 19
Appendix C Additional Results
C.1 Ablation Studies
Table 8: Ablation on different components. We report the average performance on RULER 300k and BABILong 512k datasets using Mistral-NeMo-Instruct-2407 model. The best values are highlighted in bold.
@@ C.1 Ablation Studies
   w/ Bad heads 44.7 22.8
@@
In this section, we conduct an ablation study to evaluate the key components contributing to the effectiveness of our approach. Specifically, we analyze the impact of (1) the choice of the recurrent compression method and (2) the selection of attention heads used for retrieval. We assess model performance on the BABILong and RULER benchmarks at context lengths of 512k and 300k, respectively.
@@
Choice of recurrent compression method. To demonstrate the generality of our approach, we replace our recurrent compression component with alternative methods, namely StreamingLLM and TOVA. While H2O yields the best results, Table 8 shows that other compression methods achieve comparable performance. This further highlights the flexibility of our framework and its potential for even higher performance with more advanced compression techniques.
@@
Choice of attention heads. To examine the importance of attention head selection for embedding construction, we replace the selected heads with (1) four randomly chosen heads and (2) four suboptimal ”bad” heads, identified based on poor performance on both synthetic datasets used for head selection. As shown in Table 8, the heads selected by our mechanism achieve the best performance, demonstrating its effectiveness. Random heads generally show lower but reasonable performance. In contrast, using bad heads results in a substantial performance drop on both benchmarks, underscoring the importance of proper attention head selection to ensure effective embedding construction.
C.2 Embedding Head Identification for Pattern Matching Task
@@
In this section, we present the distribution of MNR scores measured with our pattern matching dataset, similarly to what we presented in Table 1 and Figure 2. The corresponding results for pattern matching dataset is presented in Table 9 and Figure 4. The retrieval performance of QKV heads often outperform that of the hidden states, similarly to the case of multi-hop QA datasets.

### v2

@@ 4.5 Ablation Studies
Table 6: Ablation study and efficiency analysis. (a) We report the average performance on RULER 300k and BABILong 512k datasets using Mistral-NeMo-Instruct-2407 model. (b) We compare the inference time and peak memory usage required for generating 10 tokens conditioned on 256k inputs. All measurements are made with the Mistral-NeMo-Instruct-2407 model on a single H100 GPU, and are averaged over 10 samples. The best values are highlighted in bold.
(a) Ablation study.
@@
     w/ Worst heads 44.7 22.8
     w/ Kernel size 5 18.4 36.8
     w/ Kernel size 17 39.4 45.8
(b) Efficiency Analysis.
Time (s) Memory (GB)
StreamingLLM 36.58 37.34
H2O 41.33 37.85
TOVA 39.46 37.06
InfiniPot 40.90 37.06
InfLLM 129.14 51.62
REFORM (Ours) 27.24 35.00
@@
We conduct an ablation study to evaluate the key components contributing to the effectiveness of our approach. Specifically, we analyze the impact of (1) the choice of the recurrent compression method, (2) the selection of attention heads used for retrieval, and (3) size of the maxpool kernel applied during the gather stage. We assess model performance on the BABILong and RULER benchmarks at context lengths of 512k and 300k, respectively.
@@
Choice of recurrent compression method. To demonstrate the generality of our approach, we replace our recurrent compression component with alternative methods, namely StreamingLLM and TOVA. While H2O yields the best results, Table˜6(a) shows that other compression methods achieve comparable performance. This further highlights the flexibility of our framework and its potential for even higher performance with more advanced compression techniques.
@@
Choice of attention heads. To examine the importance of attention head selection for embedding construction, we replace the selected heads with (1) four randomly chosen heads and (2) four worst heads, identified based on poor performance on both synthetic datasets used for head selection. As shown in Table˜6(a), the heads selected by our mechanism achieve the best performance, demonstrating its effectiveness. Random heads generally show lower but reasonable performance. In contrast, using bad heads results in a substantial performance drop on both benchmarks, underscoring the importance of proper attention head selection to ensure effective embedding construction.
Pooling kernel size. In the on-demand cache recomputation phase, we apply max-pooling over 129-token windows to smooth token-level similarity scores. As demonstrated in Table˜6(a), reducing the pooling size to 5 or 17 tokens significantly degrades performance, highlighting the importance of pooling to maintain contextual information during the recomputation.
4.6 Efficiency Analysis
To highlight the efficiency benefits of our approach, we measure the peak memory usage and inference time required for processing a long input. We outline the results in Table˜6(b). InfLLM suffers from high inference time due to frequent memory transfer between CPU and GPU and requires large memory to store the cache. Recurrent methods offer faster inference at lower memory costs, enjoying the benefits of using a fixed-size KV cache. Our approach shows even lower latency and memory requirements compared to the recurrent baselines thanks to the early exit, which saves computation as well as memory by removing the need to keep the KV cache for the upper layers. We present further latency breakdown analysis in Section˜D.2, highlighting that REFORM keeps both time to first token (TTFT) and time per output token (TPOT) minimal.
5 Conclusion
We introduce REFORM, a novel inference framework for efficient long-context processing. REFORM incrementally processes input chunks while maintaining a compressed KV cache, extracting key QKV states to construct cross-layer context embeddings. An early-exit strategy enhances efficiency, and a similarity-based selection mechanism identifies and gathers essential tokens for KV cache recomputation. REFORM outperforms existing methods across long-context benchmarks while reducing inference time and memory usage. Furthermore, its modality-agnostic design makes it applicable to a wide range of use cases including multi-modal applications.
Limitations and Future work. One limitation is that the effectiveness of REFORM depends in part on the gather stage’s ability to identify informative tokens. While our results across diverse benchmarks demonstrate that this step is generally reliable, edge cases may exist where the model’s selection could be further optimized. Another limitation is the redundant attention score computation in the current implementation. Our current implementation recomputes attention scores using matrix multiplication to compute the eviction scores, as Flash Attention does not output attention weights. Integrating eviction score computation into the Flash Attention kernel could further improve REFORM’s efficiency.
Also, we directly adopt the token eviction criteria proposed by H2O (zhang2023h2o) as the compression component of our framework. As a future direction, we plan to investigate more sophisticated compression approaches tailored for constructing the context embeddings. Furthermore, applications to more diverse data modalities such as audio and video is another promising direction to explore.
Acknowledgements
For WS and JS: This work was supported by Institute for Information & communications Technology Promotion(IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II190075 Artificial Intelligence Graduate School Program (KAIST); No. RS-2024-00509279, Global AI Frontier Lab; RS-2022-II220959, Few-shot Learning of Causal Inference in Vision and Language for Decision Making).
WS expresses his sincere gratitude to Jihoon Tack for his insightful discussions, support, and mentorship throughout the course of this work.
Appendix A REFORM algorithm
We illustrate the overall process of REFORM through a pseudocode in Algorithm˜1.
Algorithm 1 Overview of REFORM
procedure ForwardChunk($\text{chunk},\text{cache},\text{emb}$)
   /* Initialize hidden states */
  $\text{hs}\leftarrow\text{input}$
   /* Forward with early exit */
  for layer in model_layers[:early_exit_layer] do
   $\text{hs},\text{cache},\text{qkv}\leftarrow\text{layer.Forward}(\text{hs},\text{cache})$
    /* Save selected embeddings */
   $\text{emb.SaveSelected}(\text{qkv})$
  end for
   /* Evict less important tokens */
  $\text{cache}\leftarrow\text{Compress}(\text{cache})$
  return $\text{cache},\text{emb}$
end procedure
procedure REFORM(input)
   /* Initialize */
  $\text{cache},\text{emb}\leftarrow\text{EmptyInit()}$
   /* Prepare input chunks */
  $\text{context},\text{query}\leftarrow\text{SplitQuery}(\text{input})$
  $\text{chunks}\leftarrow\text{ChunkInputs(context) + [query]}$
   /* Recurrent chunked forwarding */
  for $c_{i}$ in chunks do
   $\text{cache},\text{emb}\leftarrow\text{ForwardChunk}(c_{i},\text{cache},\text{emb})$
  end for
   /* Gather relevant inputs */
  $\text{relevant\_inputs}\leftarrow\text{GatherRelevant(input, emb)}$
   /* On-demand recomputation */
  $\text{cache}\leftarrow\text{model.Forward}(\text{relevant\_inputs})$
  return cache
end procedure
Appendix B Extended related works
Extending LLMs to handle extremely long inputs. To extend the context windows of Large Language Modles (LLMs) efficiently, various approaches have been proposed. A significant body of work focuses on modifying positional embeddings. These include scaling Rotary Positional Embeddings (RoPE) [su2024roformer] beyond the model’s context limit [chen2023extending, kaiokendev2023linear, bloc972023ntk, peng2023yarn], applying attention masks [han2023lm], or adjusting the relative distances between tokens to fall within a predefined range [rerope2023, jin2024llm]. Another line of research explores fine-tuning techniques to adapt models for longer contexts [zhu2023pose, chen2023longlora]. While these methods enable models to handle extended inputs, they do not address the significant computational and memory costs introduced by the self-attention mechanism, limiting their practical utility for extremely long contexts. Hence, we did not include them as baselines in our experiments.
Other approaches for efficient long context processing. Together with the recurrent KV cache compression approaches, a large volume of recent works focus on reducing the size of the KV cache to enable more efficient inference at long contexts. For example, SnapKV [li2024snapkv] proposes to forward the full input through the model, and then compress the cache by evicting tokens based on attention scores. While efficient at decoding-time, it requires the model to first process the full input, and therefore is not applicable to extremely long inputs that exceed the model’s pre-trained context window. Alternatively, HOMER [song2024hierarchical] proposes to use a hierarchical divide-and-conquer approach to combine the encoding and eviction process. Some works propose to further enhance KV cache compression by merging tokens instead of evicting them [wang2024model, dong2024get], but their experiments also only consider inputs within the model’s context limit, and their extrapolation capabilities remain unknown. Some recent works propose another direction to keep the full cache only for some selected attention heads known as ‘retrieval heads’ [tang2024razorattention, xiao2024duoattention], reducing the memory burden of preserving the full KV cache. Other works investigate quantization [hooper2024kvquant, kang2024gear] and low-rank cache compression [singhania2024loki] to further reduce the memory requirements of the KV cache. However, these methods also cannot extrapolate to longer sequences beyond the model’s pre-trained context limit.
Appendix C Experimental Details
C.1 Evaluating Retrieval Heads and Embeddings
Dataset preparation. To evaluate the embeddings, we constructed a synthetic dataset based on multi-hop question answering. In this setup, we embedded documents from the HotPotQA dataset [yang2018hotpotqa] at random positions within a long text corpus derived from the WikiText dataset [merity2016pointer]. Each question was appended at the end of the context, and token-level labels were created, where tokens from the golden documents were marked as ground truth. All samples were designed to be 8k tokens long, which is within the context window of the Mistral-7B-Instruct-v0.2 model.
Embedding extraction. To simulate long-context scenarios where full attention computation is infeasible due to computational or memory constraints, we employed a recurrent chunk forwarding method based on H2O [zhang2023h2o], elaborated in Section˜3.1. For attention, we compute the retrieval scores using the dot product between query states (Q) and the key states (K) without applying positional encoding. For all other embeddings, we compute the significance scores using cosine similarity between question embeddings and context embeddings, followed by max-pooling over question tokens. Additionally, retrieval scores for each context token were smoothed by mean-pooling with 20 neighboring tokens.
C.2 Detailed Head Selection Process
Step 1: Data preparation. We construct two synthetic tasks for attention head selection: a pattern matching task and a multi-hop question answering task. For the pattern matching task, we insert sentences into the WikiText corpus with the format: “The value corresponding to the id key is value.” where both the key and value are random 10-character ASCII strings. For the multi-hop QA task, we use passages from HotPotQA that contain the information required to answer a given question. The tokens corresponding to this key information are labeled as “golden tokens.”
Step 2: Context encoding. We apply recurrent chunked forwarding using H2O to each sample in the dataset, extracting token-level QKV embeddings from every layer and attention head.
Step 3: Significance score computation. For each QKV head in each layer, we compute a token-wise significance score. This is done by (1) calculating cosine similarity between context tokens and question tokens, (2) aggregating the similarity scores over the question tokens via max-pooling, and (3) smoothing the scores using mean pooling across a 20-token window.
Step 4: Performance measurement and head selection. We evaluate the effectiveness of each attention head using the Mean Normalized Rank (MNR) score, as described in Section˜C.1. Based on these scores, we select four heads: two that perform best on the pattern matching task, and two that perform best on the multi-hop QA task.
C.3 Multimodal Evaluations
C.4 Comparison against RAG
Experiment setup. We follow the setup in OP-RAG [yu2024defense] for the RAG experiments, segmenting the inputs to 128-token chunks and preserving the order of the chunks instead of rearranging them according to the retrieval scores. We use Mistral-NeMo-Instruct-2407 as the base LLM. We use BM25 [robertson1995okapi] as the sparse retriever, and bge-large-en-v1.5 [bge_embedding] as the dense retriever. For each sample, 8k tokens are retrieved in total, matching the KV size with our approach to ensure fair comparison.
C.5 Efficiency Measurements
Experiment setup. We measure the average inference time and peak memory usage for generating 10 tokens conditioned on 256k tokens. All measurements are made on a single H100 GPU, and we apply Flash Attention 2 [dao2023flashattention2] for all measurements. We further elaborate the experiment setup for InfLLM, as the inference speed and memory consumption can largely vary depending on the configuration. We use the default configuration provided in their GitHub repository, while modifying the number of retrieved blocks to keep 32k active tokens in the cache. The maximum number of blocks cached in GPU was set to be the twice as large as the number of retrieved blocks, following the convention in their official configuration file.
C.6 Embedding Construction and Similarity Search for REFORM
Embedding head selection. We construct the context embeddings by combining four QKV embeddings, where two heads are identified using the pattern matching dataset and the other two are identified using the multi-hop QA dataset. To balance between performance and efficiency gains, we select the top-performing heads from layers with depth under 70% for pattern matching heads. See Section˜D.1 for a more detailed discussion.
For Mistral-NeMo-Instruct-2407, the following heads are used:
  1. 1.
    Query head 9 at layer 15
  2. 2.
    Value head 5 at layer 19
  3. 3.
    Value head 0 at layer 27
  4. 4.
    Value head 7 at layer 27
For Qwen2.5-7B-Instruct, the following heads are used:
  1. 1.
    Value head 3 at layer 7
  2. 2.
    Key head 0 at layer 14
  3. 3.
    Value head 3 at layer 14
  4. 4.
    Value head 0 at layer 19
For Qwen2.5-Coder-1.5B-Instruct, the following heads are used:
  1. 1.
    Query head 3 at layer 8
  2. 2.
    Value head 1 at layer 11
  3. 3.
    Key head 0 at layer 14
  4. 4.
    Value head 0 at layer 15
For Qwen2.5-Coder-7B-Instruct, the following heads are used:
  1. 1.
    Value head 2 at layer 13
  2. 2.
    Key head 0 at layer 14
  3. 3.
    Value head 3 at layer 14
  4. 4.
    Query head 4 at layer 14
For Pixtral-12B-2409, the following heads are used:
  1. 1.
    Value head 3 at layer 10
  2. 2.
    Value head 5 at layer 19
  3. 3.
    Value head 0 at layer 27
  4. 4.
    Value head 7 at layer 27
Appendix D Additional Results
D.1 Embedding Head Identification for Pattern Matching Task
@@
In this section, we present the distribution of MNR scores measured with our pattern matching dataset, similarly to what we presented in Table˜1 and Figure˜2. The corresponding results for pattern matching dataset is presented in Table˜7 and Figure˜4. The retrieval performance of QKV heads often outperform that of the hidden states, similarly to the case of multi-hop QA datasets.
@@
D.2 Latency Breakdown
Here, we extend our latency analysis to three context lengths (only 256k in the main text), and separately present the pre-fill and decoding latency in Table˜8(a) and Table˜8(b).
In Table˜8(a), InfiniPot and InfLLM shows slower processing speed compared to StreamingLLM due to the additional computation associated with dynamic cache eviction and retrieval. On the other hand, REFORM consistently shows faster prefill time compared to all baselines, thanks to the early-exit optimization. Further latency breakdown of the Compress+Gather stage and Recompute stage suggests that the recomputation overhead is minimal, taking under 1.52% of the total prefill time.
Table˜8(b) shows a similar trend for the decoding time. InfiniPot and InfLLM shows slower decoding time compared to StreamingLLM due to the additional operations. On the other hand, REFORM also shows improved decoding speed as the generation is conditioned on KV cache created with a small recomputation budget (8k), unlike the baselines that condition on the full compressed cache (32k). (Note that these budgets are the hyperparameters used for the Mistral-Nemo experiments in our main text.)
Table 8: Latency Breakdown. (a) Time to first token (seconds) measurements and (b) time per output token (seconds) measurements (Mistral-Nemo-Instruct-2407, single H200, averaged over 20 runs and 200 tokens generated per measurement).
(a) Time to first token (seconds).
Model 256k 512k 1M
StreamingLLM 30.59 x68.22 143.57
InfiniPot 35.77 x73.67 149.64
InfLLM 95.71 213.23 474.96
REFORM (Ours) 26.24 x53.68 108.64
     - Compress + Gather 25.84 x53.28 108.24
     - Recompute x0.40 xx0.40 xx0.40
(b) Time per output token (seconds)
Model 256k 512k 1M
StreamingLLM 0.111 0.111 0.111
InfiniPot 0.256 0.256 0.256
InfLLM 0.259 0.267 0.329
REFORM (Ours) 0.040 0.040 0.040
D.3 Evaluation with a Larger Model
In this section, we present additional results using the Qwen2.5-32B-Instruct model on the RULER and BABILong benchmarks. As shown in Table˜9, REFORM consistently outperforms both H2O and InfiniPot across all evaluation settings, demonstrating the scalability and robustness of our method on larger models.
Table 9: Performance of Qwen2.5-32B-Instruct. We report the performance on key long-context benchmarks for H2O, InfiniPot, and REFORM. The best values are highlighted in bold.
RULER RULER RULER RULER BABILong
Single 300k Multikey 300k Multivalue 300k Multiquery 300k 256k
H2O 38.7 x2.7 14.0 x6.0 31.0
InfiniPot 69.3 19.3 40.0 74.0 54.2
REFORM (Ours) 100.0 90.0 96.0 100.0 67.6
Appendix E Further Discussions
E.1 Complexity Analysis
Time and memory complexity. Recurrent baselines such as StreamingLLM have a time complexity of $O(L)$, where $L$ is the input length, assuming a fixed (or bounded) chunk size and query length. Each recurrence step involves a constant amount of computation, and the total number of steps scales linearly with the input length. Their memory complexity is $O(1)$, since only a fixed-size KV cache is maintained throughout.
In contrast, random-access baselines like InfLLM incur a time complexity of $O(L^{2})$, primarily due to periodic cache lookups across the full input length. These methods also require $O(L)$ memory, as they store the entire KV cache in memory. This memory burden is significant, often requiring CPU offloading, which further increases latency.
REFORM maintains a time complexity of $O(L)$ through its recurrent chunked processing. Although it has an $O(L)$ memory cost due to storing token-level embeddings, these embeddings are very small in practice. Moreover, the early-exit strategy significantly reduces memory and compute requirements. Consequently, REFORM achieves faster speed and lower memory usage than both baselines (as demonstrated in Table˜6(b)), highlighting the practical efficiency of our method.
Amount of offline computation. While REFORM involves offline computation for head selection, the computational overhead of head selection is minimal in practice. The selection is performed only once using synthetic inputs of 8k tokens, and the chosen heads are reused across all inference runs, independent of the downstream task or input domain. By selecting heads using short inputs and reusing them across long-context inference, REFORM keeps the additional computation minimal. Quantitatively, the entire head selection procedure requires roughly the same amount of computation (in FLOPs) as processing just two 1M-token inputs with H2O.
E.2 Theoretical Analysis on Head Selection
Our attention head selection is motivated by findings from mechanistic interpretability literature, which show that different heads specialize in different functions [voita2019analyzing, ren2024identifying]. Some heads are more useful than others for tasks like pattern recognition or reasoning. We leverage this diversity by selecting only a few high-performing heads, both for better identification performance and to keep the size of token-level embeddings tractable (including all heads would be memory-intensive). Our method thus aims for functional specialization as well as computational efficiency.
Although we use cosine similarity rather than attention weights to evaluate heads, our approach shares key insights with existing interpretability work. Prior studies show that lower-layer heads often detect syntactic or structural features, mid-layer heads handle semantic reasoning, and upper-layer heads guide output generation [zheng2024attention]. This aligns with our empirical findings: the most useful heads for pattern matching tend to appear in lower and upper layers, while those for multi-hop QA (i.e. semantic understanding) concentrate in the mid-layers. These trends are reflected in our MNR visualizations (Figure˜2 and Figure˜4). Also, Skean et. al. [skean2025layer] suggests that the middle layer representations are more useful than the final layer representations for 32 MTEB (Massive Text Embedding Benchmark, [muennighoff2022mteb]) tasks, similar to what we observe for the multi-hop QA task.
Furthermore, our choice of the head identification tasks is rooted in information retrieval literature, which shows that hybrid retrievers [bruch2023analysis, chen2022out] combining sparse (structural) and dense (semantic) retrievers outperforms either alone. Inspired by this, we use two distinct head evaluation tasks: pattern matching and multi-hop QA. The former identifies heads that capture structural similarity, while the latter targets heads that encode semantic relevance. This dual-task strategy helps us select a complementary set of heads that are robust across diverse long-context scenarios.
Appendix F Broader Impacts
We believe that the high capability and flexibility will aid everyday use of large foundation models, by extending the model capabilities to efficiently and effectively handle very long contexts. On the other hand, such capabilities of REFORM could potentially enable malicious parties to analyze vast amount of data, enhancing the capabilities of autonomous systems that could be used for manipulation or misinformation.
Appendix G License Information for Datasets and Models
Here, we provide the license for all datasets and models used in our experiments. Apache 2.0 license is applied for Babilong, RULER, Mistral-Nemo-Instruct-2407, Qwen2.5-Coder family, and Pixtral-12B-2409. BSD license is also applied for some parts of Babilong dataset. MIT license is applied to Needle-in-a-Haystack, InfiniteBench, RepoEval, WikiText, and bge-large-en-v1.5. CC BY-SA 4.0 is applied for HotPotQA.